Session: Jun 2014 (session 2)
Wavelength: 18 cm
To give feedback on this experiment, click
(C. Nocita at Thu May 29 19:43:49 2014) Success:
(bo xia at Fri May 30 00:25:13 2014) Success:
(Geert Kuper at Fri May 30 07:49:50 2014) Minor failures: RECO
Tadumax was out of sync. (tickerror). Needed reboot and resync.
Missed from 13:44 UT - 14:03 UT.
(Gawronski at Fri May 30 08:24:53 2014) Minor failures:
Missed from 14:36 UT - 15:00 UT - DBBC was rebooted. Swapped bitstreams in our data
Svetloe (Slobojaninova N. at Fri May 30 11:27:42 2014) Success:
Zelenchukskaya (Kiselev P. at Fri May 30 11:28:31 2014) Success:
Scans from No0107 to No0127 missed due to participation in intensive geo session
Badary (Naumov, Podlosinskaya at Fri May 30 11:29:39 2014) Success:
Pause for intensive geo session (scans no0107-0127).
(Jun Yang at Mon Jun 02 09:09:11 2014) Success:
No known problems.
(Jimmy Cullen at Tue Jun 03 09:30:23 2014) Success:
No known problems
(Jimmy Cullen at Wed Jun 04 11:47:02 2014) Success:
Due to a bogie failure, the Lovell was replaced by the MkII for
(Uwe Bach at Thu Jun 05 13:13:39 2014) Success: